OpenStudy (anonymous):

write an equation of a line parallel to y=3x+7 and passes through the point (0,-1)

OpenStudy (anonymous):

notice is says "write" not "find" because you do this with your eyeballs, not algebra what is the slope of your line?

OpenStudy (anonymous):

i mean what is the slope of \(y=3x+7\) ?

OpenStudy (anonymous):

for the line of the form y = mx + c (y= 3x + 7) slope is given by m use : parallel lines have equal slopes and the eq of the line passing through (x1,y1) is given by y-y1 = m (x-x1) here the reqd line passes through (0,-1) and

OpenStudy (anonymous):

nah just write it